Goods Market Money Market
C = 250 + 0.8YD L = 0.25Y – 62.5i
YD = Y + TR – T Ms/P = 250
T = 100 + 0.25Y
I = 300 – 50i
G = 350; TR = 150

Goods market equilibrium condition: Y = C + I + G + X-M
Money market equilibrium condition: L = Ms/P

a) What is the equation that describes the IS curve (YIS)?
1
Expert's answer
2019-10-09T11:06:39-0400

a) The equation that describes the IS curve (YIS) is:

Y = C + I + G = 250 + 0.8(Y + TR – T) + 300 – 50i + 350 = 900 + 0.8(Y + TR – T) – 50i.